Find the total area of four walls in each of the rectangular rooms having the following dimensions:
Length=10m,breadth=5m,height=4m
Answer is 120
Plz tell me how I tried it but failed

Answer by jim_thompson5910(35256)

One wall's area = (length)*(height) = 10*4 = 40 square meters

The parallel wall on the opposite side is also 40 square meters
----------------------------------------------------------------

Other wall's area = (breadth)*(height) = 5*4 = 20 square meters

The parallel wall on the opposite side is also 20 square meters
----------------------------------------------------------------

Add the areas: 40+40+20+20 = 120

So that's why the answer is 120 square meters